You want to construct a investment portfolio consisting of $400,000 in Treasury Bills yielding 3% and $600,000 in a Market Portfolio with an expected market return of 10%.

            a) What is the market risk premium?

            b) What is the portfolio’s risk premium?
Answer #1

HI

Market risk premium is the difference between market return and risk free rate

Here risk free rate = T bill rate = 3%

market return = 10%

So a) market risk premium = 10- 3 = 7%

b) Portfolio risk premium is the difference between portfolio return and risk free rate

Here Portfolio return = 400,000/(400,000+600,000)*3% + 600,000/(400,000+600,000)*10%

= 0.4*3% + 0.6*10%

=1.2 + 6 = 7.2%

So Portfolio risk premium = 7.2 - 3 = 4.2%

Thanks